OpenStudy (anonymous):

use the intermediate value theorem to prove that p(x)=X^7+x^6+x^5+x^4+x^3+x^2+x+1 has at least one zero

OpenStudy (anonymous):

Try plugging in values like \(x=0\) and \(x=-2\). You won't find a positive root to this, because \(p(\text{any positive})=\text{positive}\). At \(x=0\), you have \(p(0)=0+0+0+0+0+0+0+1=1\), at \(x=-1\), you have \(p(-1)=-1+1-1+1-1+1-1+1=0\), and so on. To apply the IVT, you need only show that \(p(x)\) attains a negative number.